OpenStudy (anonymous):

Find the slope of segment CD with endpoints C(−5, 4) and D(2, 1)

pooja195 (pooja195):

Hi there :) Simply use the slope formula \[ \huge~\rm~\bf\color{blue}{ \frac{y_2-y_1}{x_2-x_1}=slope} \]

pooja195 (pooja195):

Need help?

OpenStudy (anonymous):

1-4/2-(-5)= -3/7 ?

pooja195 (pooja195):

Perfect :)

OpenStudy (anonymous):

:)